OpenStudy (anonymous):

how do I solve this problem? Find the domain of the following: f(x) = x over x-squared - 3x - 10.

OpenStudy (anonymous):

the denominator is a quadratic right? X^2-3x-10

OpenStudy (anonymous):

yes, it is

OpenStudy (anonymous):

so set it to = 0 and apply the quadratic formula

OpenStudy (anonymous):

so for a=1 b=-3 and c=-10 ?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

i got 3/2 +or -7/2

OpenStudy (anonymous):

wow, i haven't answered it yet, but thank you for your answer. I have on more question. How to i solve f(x) = -2 over square root of 4+x ?

OpenStudy (anonymous):

so for the 1st one its X cant be 5 or 2 got it?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i meant thats the answer to the first problems. try to work it out and see if you understand it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[-2/\sqrt{4+x}\]

OpenStudy (anonymous):

is that the second problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you cant have a negative square root so set the squareroot of \[\sqrt{4+x}\] greater or equal to zero

OpenStudy (anonymous):

i mean 1

OpenStudy (anonymous):

is the answer -3? for problem 2?

OpenStudy (anonymous):

yup! it means that x have to be greater or equal to -3

OpenStudy (anonymous):

wow, thank you :)

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

yes, problem two does. I tried to solve problem 1 and i got \[x = 3/2 \pm \sqrt{151} / 2 i\]

OpenStudy (anonymous):

do i deserve a medal? =)

OpenStudy (anonymous):

:) how do i give you a medal?

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4ac}\] all over 2a

OpenStudy (anonymous):

click good answer

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you might plug something wrong

OpenStudy (anonymous):

i approximated the answer to find what x equals and i got x = 7.64 and x = -4.64. is that what you meant by 7 and -5?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

ok so b = -3 a = 1 c = -10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3\pm \sqrt{9-4(1)(-10)}\]

OpenStudy (anonymous):

sorry i donw know how to do the all over sign

OpenStudy (anonymous):

ok. it's over 2 right?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

\[3\pm \sqrt{49}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

i see what i did wrong

OpenStudy (anonymous):

\[3/2\pm7/2\] sqrt of 49 = 7

OpenStudy (anonymous):

yes, i got it. Thank you :) I have to go now.

OpenStudy (anonymous):

welcome